I am conducting a research using six areas with a population of 169,894. i used the Yamane sampli...

i am conducting a research using six areas with a population of 169,894. i used the Yamane sampling technique to get a sample of 400. Yhe problem here is i dont have the population for the respective areas so how do i determine the sample size for these six areas which will be a general representative of the population to administer my questionnaires.
Answer #1

The answer to this question will be quite subjective and may change person to person and technique to technique.

Following is my answer,

To decide the sample size for these size areas you can consider specific proportion of the sample points should come from each population from all six areas.

eg- Suppose you have 1000 population size over six different areas,

Then, let p = 100/6 = 16.66 = 17%

And we want sample to choose of size 500 including all areas.

So, from each area will choose around 16 or 17 % observation of my sample size.

Hence, 500*0.16 = 80

Will choose, 80 observations from each area to conclude the complete sample of size 500 even though i dont know the sample size for each area.